# AfriMTE and AfriCOMET: Enhancing COMET to Embrace Under-resourced African Languages

### v2

@@ Abstract
Jiayi Wang, David Ifeoluwa Adelani, Sweta Agrawal††thanks: The authors contribute equally to this work and are considered co-third authors. , Marek Masiak††thanks: Currently at the University of Oxford, UK. , Ricardo Rei, Eleftheria Briakou, Marine Carpuat, Xuanli He, Sofia Bourhim, Andiswa Bukula, Muhidin Mohamed, Temitayo Olatoye, Tosin Adewumi, Hamam Mokayede, Christine Mwase, Wangui Kimotho, Foutse Yuehgoh, Anuoluwapo Aremu, Jessica Ojo, Shamsuddeen Hassan Muhammad, Salomey Osei, Abdul-Hakeem Omotayo, Chiamaka Chukwuneke, Perez Ogayo, Oumaima Hourrane, Salma El Anigri, Lolwethu Ndolela, Thabiso Mangwana, Shafie Abdi Mohamed, Ayinde Hassan, Oluwabusayo Olufunke Awoyomi, Lama Alkhaled, Sana Al-Azzawi, Naome A. Etori, Millicent Ochieng, Clemencia Siro, Samuel Njoroge, Eric Muchiri, Wangari Kimotho, Lyse Naomi Wamba Momo, Daud Abolade, Simbiat Ajao, Iyanuoluwa Shode, Ricky Macharm, Ruqayya Nasir Iro, Saheed S. Abdullahi, Stephen E. Moore, Bernard Opoku, Zainab Akinjobi, Abeeb Afolabi, Nnaemeka Obiefuna, Onyekachi Raphael Ogbu, Sam Brian, Verrah Akinyi Otiende, Chinedu Emmanuel Mbonu, Sakayo Toadoum Sari, Yao Lu, Pontus Stenetorp University College London, UK, Masakhane NLP, University of Maryland, USA, Unbabel, Instituto Superior Técnico, INESC-ID, ENSIAS, Morocco, SADiLaR, South Africa, Aston University, UK, University of Eastern Finland, Finland, Luleå University of Technology, Sweden, Fudan University, China, Conservatoire National des Arts et Métiers, France, Lelapa AI, South Africa, Imperial College London, UK, University of Deusto, Spain, University of California, USA, Lancaster University, UK, Mohammed V University, Morocco, Jamhuriya University Of Science and Technology, Somalia, LAUTECH, Nigeria, The College of Saint Rose, USA, University of Minnesota -Twin Cities, USA, Microsoft Africa Research Institute, University of Amsterdam, Netherlands, The Technical University of Kenya, AIMS, Cameroon, KU Leuven, Belgium, HausaNLP, UCAS, China, Kaduna State University, Nigeria, University of Cape Coast, Ghana, Ghana NLP, Kwame Nkrumah University of Science and Technology, Ghana, New Mexico State University, USA, USIU-Africa, UNIZIK, Nigeria, AIMS, Senegal Corresponding emails: {ucabj45,d.adelani}@ucl.ac.uk, p.stenetorp@cs.ucl.ac.uk
@@ Acknowledgments
David Adelani acknowledges the support of DeepMind Academic Fellowship programme. Ricardo Rei is supported by the European Union’s Horizon Europe Research and Innovation Actions (UTTER: contract 101070631) and by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI). Pontus Stenetorp would like to acknowledge the helpful proofing feedback from several viewers while finalizing the submission. We are grateful to Prof. Antonios Anastasopoulos from George Mason University for releasing the “WMT African” dataset for our experiments. This work is supported in part by Oracle Cloud credits and related resources provided by Oracle. Finally, we are grateful to OpenAI for providing Masakhane with API credits through their Researcher Access API program for the evaluation of the GPT-4 large language models.

### v3

@@ Abstract
Jiayi Wang1, David Ifeoluwa Adelani1,2, Sweta Agrawal3,††thanks: The authors contribute equally to this work and are considered co-third authors. , Marek Masiak1,∗,††thanks: Currently at the University of Oxford, UK. , Ricardo Rei4,5,6, Eleftheria Briakou3, Marine Carpuat3, Xuanli He1, Sofia Bourhim7, Andiswa Bukula8, Muhidin Mohamed9, Temitayo Olatoye10, Tosin Adewumi11, Hamam Mokayed11, Christine Mwase12, Wangui Kimotho2, Foutse Yuehgoh13, Anuoluwapo Aremu2, Jessica Ojo14,2, Shamsuddeen Hassan Muhammad15,2,29, Salomey Osei16,2, Abdul-Hakeem Omotayo17,2, Chiamaka Chukwuneke18,2, Perez Ogayo2, Oumaima Hourrane2, Salma El Anigri19, Lolwethu Ndolela2, Thabiso Mangwana2, Shafie Abdi Mohamed20, Ayinde Hassan21, Oluwabusayo Olufunke Awoyomi22, Lama Alkhaled11, Sana Al-Azzawi11, Naome A. Etori23, Millicent Ochieng24, Clemencia Siro25, Samuel Njoroge26, Eric Muchiri2, Wangari Kimotho27, Lyse Naomi Wamba Momo28, Daud Abolade2, Simbiat Ajao2, Iyanuoluwa Shode2, Ricky Macharm2, Ruqayya Nasir Iro29, Saheed S. Abdullahi30,31, Stephen E. Moore32,33, Bernard Opoku34,2, Zainab Akinjobi35,2, Abeeb Afolabi2, Nnaemeka Obiefuna2, Onyekachi Raphael Ogbu2, Sam Brian2, Verrah Akinyi Otiende36, Chinedu Emmanuel Mbonu37, Sakayo Toadoum Sari38, Yao Lu1, Pontus Stenetorp1 1University College London, UK, 2Masakhane NLP, 3University of Maryland, USA, 4Unbabel, 5Instituto Superior Técnico, 6INESC-ID, 7ENSIAS, Morocco, 8SADiLaR, South Africa, 9Aston University, UK, 10University of Eastern Finland, Finland, 11Luleå University of Technology, Sweden, 12Fudan University, China, 13Conservatoire National des Arts et Métiers, France, 14Lelapa AI, South Africa, 15Imperial College London, UK, 16University of Deusto, Spain, 17University of California, USA, 18Lancaster University, UK, 19Mohammed V University, Morocco, 20Jamhuriya University Of Science and Technology, Somalia, 21LAUTECH, Nigeria, 22The College of Saint Rose, USA, 23University of Minnesota -Twin Cities, USA, 24Microsoft Africa Research Institute, 25University of Amsterdam, Netherlands, 26The Technical University of Kenya, 27AIMS, Cameroon, 28KU Leuven, Belgium, 29HausaNLP, 30SIAT-CAS, China, 31Kaduna State University, Nigeria, 32University of Cape Coast, Ghana, 33Ghana NLP, 34Kwame Nkrumah University of Science and Technology, Ghana, 35New Mexico State University, USA, 36USIU-Africa, 37UNIZIK, Nigeria, 38AIMS, Senegal Corresponding emails: {ucabj45,d.adelani}@ucl.ac.uk, p.stenetorp@cs.ucl.ac.uk
@@ Acknowledgments
David Adelani acknowledges the support of DeepMind Academic Fellowship programme. Ricardo Rei is supported by the European Union’s Horizon Europe Research and Innovation Actions (UTTER: contract 101070631) and by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI). Pontus Stenetorp would like to acknowledge the helpful proofing feedback from several viewers while finalizing the submission. We are grateful to Prof. Antonios Anastasopoulos from George Mason University for releasing the “WMT African” dataset for our experiments. This work is supported in part by Oracle Cloud credits and related resources provided by Oracle. This work is also supported in part by Microsoft Research via their Accelerate Foundation Models Research Grant. Finally, we are grateful to OpenAI for providing Masakhane with API credits through their Researcher Access API program for the evaluation of the GPT-4 large language models.